OpenStudy (anonymous):

How would you write the following into an equation? I have some lemons. Jill has two more limes than five times the number of lemons that I have. Emma has one-third as many limes as Jill. Emma has 4 limes. How many lemons do I have?

OpenStudy (anonymous):

J = 2 + 5M E = J/3 E = 4

OpenStudy (anonymous):

J for jill, m for me and e for emma